Select the value of x that is a solution for the equation.
6/x=7/x+3

Answer by waynest(281)

(6/x) = (7/x) + 3
x(6/x) = x(7/x) + 3(x)
6 = 7 + 3x
6 - 7 = 7 - 7 + 3x
-1 = 3x
-1/3 = 3x/3
(-1/3) = x
check:
6/(-1/3) = 7/(-1/3) + 3
-18 = -21 + 3
-18 = -18